OpenStudy (anonymous):

I NEED HELP PLEASE! find the solutions for g(x) = x^2 +6x + 1

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

yes please!

jimthompson5910 (jim_thompson5910):

Have you learned about the quadratic formula at all?

OpenStudy (anonymous):

yeah i just forget how to get the solutions of this formula i know i need an a, b and c after that i kind forget

jimthompson5910 (jim_thompson5910):

x^2 +6x + 1 really is 1x^2 +6x + 1

jimthompson5910 (jim_thompson5910):

1x^2 +6x + 1 is in the form ax^2 + bx + c where in this case a = 1 b = 6 c = 1

OpenStudy (anonymous):

yeah i got that far but after that I'm confused on what to do next

jimthompson5910 (jim_thompson5910):

Then you would use the quadratic formula \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

Plug in a = 1, b = 6, c = 1 \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-6 \pm \sqrt{6^2-4*1*1}}{2*1}\] \[\Large x=\frac{-6 \pm \sqrt{36-4*1*1}}{2*1}\] \[\Large x=\frac{-6 \pm \sqrt{36-4}}{2}\] I'll let you finish

OpenStudy (anonymous):

do i have to simplify the sqrt or 32?

jimthompson5910 (jim_thompson5910):

yes if you want to fully simplify the roots

OpenStudy (anonymous):

it would be -6 - 2sqrt4 over 2 and -6 + 2sqrt4 over 2

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{32} = \sqrt{16*2}\] \[\Large \sqrt{32} = \sqrt{16}*\sqrt{2}\] \[\Large \sqrt{32} = 4\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

So here is the full step by step picture. \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-6 \pm \sqrt{6^2-4*1*1}}{2*1}\] \[\Large x=\frac{-6 \pm \sqrt{36-4*1*1}}{2*1}\] \[\Large x=\frac{-6 \pm \sqrt{36-4}}{2}\] \[\Large x=\frac{-6 \pm \sqrt{32}}{2}\] \[\Large x=\frac{-6 \pm \sqrt{16*2}}{2}\] \[\Large x=\frac{-6 \pm \sqrt{16}*\sqrt{2}}{2}\] \[\Large x=\frac{-6 \pm 4\sqrt{2}}{2}\] \[\Large x=\frac{2(-3 \pm 2\sqrt{2})}{2}\] \[\Large x=\frac{\cancel{2}(-3 \pm 2\sqrt{2})}{\cancel{2}}\] \[\Large x=-3 \pm 2\sqrt{2}\]

OpenStudy (anonymous):

thank you so much!

jimthompson5910 (jim_thompson5910):

Also, the plus minus symbol \(\Large \pm\) tells us to break the solutions down so we really have these two solutions \[\Large x=-3 + 2\sqrt{2} \text{ or } x=-3 - 2\sqrt{2}\] but I find it more convenient and compact to write the two solutions as \[\Large x=-3 \pm 2\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

you're welcome